A cylindrical pipe with water flowing downward at 0.02 m3/s having top diameter of 0.08 m, bottom diameter of 0.04 m, and an elevation of 1.5 m. Determine the fluid pressure difference, kPa, across the pipe.


1
Expert's answer
2021-05-07T08:13:18-0400

"velocity 1= 0.02/(\pi(0.08/2)^2)"

"v1= 1.018*10^-5m\/s"

"velocity 2= 0.02/(\pi(0.08/2)^2)"

"v2= 2.546*10^-6m\/s"

bernoulis equation = "1\/2ev1^2 +p1+ egy1 = 1\/2ev2^2 +p1+ egy1"

"p1-p2 = 0.5ev2^2+egy2 -0.5ev1^2"

"=0.5*997.77*(2.5-10^-6)^2+ 977.77*9.8*1.5-0.5*977.77*(1.018*10^-5)^2"

="14667.219 N\/m"

= 14.667kPa